# Glyph: Scaling Context Windows via Visual-Text Compression

### v1

@@ 1 Introduction
Specially, Glyph consists of three main stages, namely, continual pre-training, LLM-driven rendering search and post-training. In the continual pre-training stage, we render large-scale long-context text into diverse visual forms, enabling the VLM to transfer its long-context capability from text tokens to visual tokens. Since the text-to-image conversion directly determines the trade-off between context compression and model performance, devising an optimal configuration of the conversion is crucial for downstream performance. To this end, we design an LLM-driven genetic search to automatically explore rendering parameters (e.g. font size, layout, resolution) to maximize compression while preserving long-context ability. The resulting configuration is then applied in the post-training stage, where we perform supervised fine-tuning and reinforcement learning to further improve the model’s performance on visualized input. An auxiliary OCR task is applied to enhance the model’s ability to recognize textual content within images, thereby better aligning its visual and textual representations, yielding the final Glyph model.
@@
  • We demonstrate that Glyph can achieve 3-4× token compression for long text sequences while preserving performance, enabling substantial improvements in memory efficiency, training and inference speed.
@@ 3.1 Overall Framework
As illustrated in Figure 2, Glyph consists of three tightly-coupled stages: (1) Continual Pre-Training, which teaches the VLM to understand and reason over rendered long texts with diverse visual styles; (2) LLM-Driven Rendering Search, automatically discovers the optimal rendering configuration for downstream tasks; and (3) Post-Training, including SFT and RL under the discovered configuration to further improve the model’s long-context capabilities. Together, these stages enable Glyph to achieve both high accuracy and significant gains in token compression, computational efficiency, and memory usage.
@@ 3.2 Task Definition
Scaling this token-based formulation to million-token contexts, however, imposes prohibitive memory and computation costs. To overcome these limitations, we reformulate the input representation through visual compression. Instead of directly feeding $\mathcal{C}$ as text tokens, we render it into a sequence of visual pages $\mathcal{V}=\{v_{1},\dots,v_{n}\}$ each containing the glyphs of multiple text segments. This allows the model to reason over a compressed but semantically equivalent input:
@@ 3.3 Continual Pre-Training
To enhance model robustness, better aligning long-text capability, we adapt diverse rendering configurations over a large amount of long-context text data. We also develop a series of rules to exclude the improper combination of rendering parameters, e.g., smaller line height than font size. Moreover, with human prior, we define several style themes, including document_style, web_style, dark_mode, code_style and artistic_pixel. These themes are designed to capture a wide range of document layouts and text styles, which can better exploit the knowledge that VLM has obtained in its pre-training stage.
@@
where $\mathcal{I}^{\!*}$ denotes an optional instruction (e.g. absent in interleaved language modeling tasks) and $\phi$ is initialized from the base VLM. This stage produces a model capable of understanding rendered text and handling long contexts, referred to as Glyph-Base.
@@ 3.5 Post Training
Table 1: Performance comparison of Glyph with leading LLMs on LongBench (%). Our model achieves competitive results in the overall average score. Best results are bolded, and second-best are underlined. Refer to Table 10 for the reset of results.
@@
Figure 3: Performance comparison of Glyph and the baseline across different context windows, demonstrating that Glyph achieves performance equivalent to longer contexts with substantially shorter context window.
@@
  • Verifiable rewards from an external LLM judge, scoring based on the accuracy of answer, which is a reference-based llm-as-a-judge with the reference being the ground truth.
@@
A persistent challenge of visual compression is the faithful recovery of fine-grained text from rendered images. Throughout both SFT and RL, we therefore incorporate an auxiliary OCR alignment task that encourages the model to correctly read and reproduce low-level textual details. The form of OCR task is the same as in the continual pre-trianing stage. In the RL stage, the reward for the OCR task is given by the Levenshtein distance.
@@ 4.1 Experimental Setup
To comprehensively evaluate the effectiveness of our method, we have conducted extensive experiments covering long-context understanding, efficiency, cross-modal generalization and several ablations and analysis. Implementation details, descriptions of baselines and benchmarks are provided in Appendix B.
@@ 4.2.1 Results on LongBench & MRCR
Tables 1 and 2 summarize overall results. Glyph achieves performance on par with or surpassing state-of-the-art text-only LLMs of similar size, including Qwen3-8B and GLM-4-9B-Chat-1M, demonstrating that Glyph remains effective on long-context tasks with large reduction in input tokens.
@@ 4.5 Ablation Study & Analysis
We compare three types of rendering configurations for SFT: (i) random sampled configuration from the pre-training sets, (ii) manually designed settings based on prior knowledge, and (iii) the configuration obtained from our search procedure. While all settings achieve comparable compression ratios, Table 5 shows that the searched configuration consistently outperforms the other two, both on average and across most individual tasks. This demonstrates the importance of systematic exploration for finding appropriate rendering strategies.
@@
To further examine the potential of our approach, we explore more aggressive compression settings. We apply a configuration with an effective $8\times$ compression ratio during post-training, and evaluate the resulting model on MRCR with sequence lengths extended from 128k to 1024k. The results (Table 7) show that Glyph successfully demonstrates the potential for $8\times$ effective context expansion, achieving performance on par with GLM-4-9B-Chat-1M and Qwen2.5-1M. This experiment highlights that our method can indeed be pushed to more extreme compression regimes while retaining performance, suggesting substantial headroom for extending usable context far beyond current limits, like a model that can deal with 4M even 8M context tokens.
@@ 5 Conclusion
Despite the effectiveness of Glyph and its strong potential for broader applications, we want to discuss several limitations of the current work that worth further exploration.
@@
Building upon the current study, several directions could further advance the proposed visual–text compression paradigm. First, rather than using a fixed rendering strategy, one promising avenue is to train adaptive rendering models that condition on the task type or user query, producing tailored visualizations that balance compression and performance. Second, enhancing the visual encoder’s capability for fine-grained text recognition and alignment with language representations could improve robustness and transferability across tasks. Third, improving the alignment between visual–text and purely textual models, for instance through knowledge distillation or cross-modal supervision, could narrow the performance gap in generalization. Fourth, our approach could be extended to broader applications, such as agent memory systems capable of managing long-term conversations or agentic contexts, and tasks that can leverage structured visual layouts for reasoning and retrieval. From the perspective of context engineering, this method offers a new way to optimize how contextual information is represented and managed. With further advances along this line, future models could go beyond the current limits of context length, effectively scaling from 1M to 10M input tokens.
@@ Appendix A Rendering Parameters
Our rendering parameters are detailed in Table 8. The best result, obtained through our LLM-driven genetic search, is presented in Fig. 6, which shows the detailed configuration and its corresponding rendering.
@@
dpi Mixture of sets: lowest (45–59), low (60–71), medium (72–119), normal ({72,80,96,100,110,120,144,150,300}), high (over 300); favor normal/medium with small prob. spikes to extremes.
@@
Table 10: The reset of results on LongBench benchmark (%), which encompasses Single-Document QA, Multi-Document QA, Summarization, Few-shot Learning, and Synthetic task.
@@ Appendix B Implementation Details
For the rendering search, we run for 5 times with 200 steps in each round, to find the optimal configuration that maximizes compression ratio while maintaining good performance.
After this, we conduct further SFT and RL training. For SFT, we train for 1.5k steps with a batch size of 32. The Adam optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.95$) is used with cosine decay and 160 warm-up steps, where the learning rate decays from 5e-6 to 2e-6. For reinforcement learning, we adopt the GRPO algorithm. Each training group samples 16 candidate responses, and degenerate samples with all-zero or all-one rewards are discarded. We apply the clip-higher trick from DAPO Yu et al. (2025b) with $\epsilon_{l}$ being 0.2 and $\epsilon_{h}$ being 0.28. Training runs for 500 iterations with a batch size of 32. We also use the Adam optimizer with constant learning rate of 1e-6.
@@
Our method rely on a strong VLM to process long-context task. Considering the impressive performance of GLM-4.1v-9B, especially in OCR and long document tasks, we choose GLM-4.1V-9B-Base as our backbone model.
@@
To conduct a comprehensive analysis of the long-context performance, we have adopt three popular benchmarks, including LongBench, MRCR and Ruler. LongBench consists of totally 21 datasets in 6 categories, covering deiverse long-context tasks. MRCR is a task proposed by , we use the OpenAI version, which consists of multi-turn conversations about writing, asking models to recall one of the context in dialogue history. Ruler is a widely-used synthetic benchmark with 11 NIAH tasks. To validate cross-model benefits, we choose the MMLongBench-Doc, which involves 130 lengthy PDF with diverse layout and images, and 1062 questions.
@@
For training, we focus on SFT since RL involves rollouts time, making it difficult to compare fairly. Moreover, running RL at very long context lengths (e.g., 64k or beyond) requires excessive memory resources. This again highlights the advantage of our approach: through compression, we can conduct RL training at 32k context length while effectively covering over 100k tokens of raw text input, where RL is prohibitively difficult for LLMs due to memory and computation demands. For SFT, we measure per-sample training time under the same number of training data using $\times8$0G H100 GPUs. For inference, we deploy both models on a single 80G H100 and measure efficiency along two axes: (i) prefill latency at batch size 1, and (ii) per-sample inference time at the maximum feasible batch size with output length set to 256 tokens. We omit the KV cache testing, because KV cache scales linearly with the sequence length, compression translates almost directly into savings of about 67% memory usage.

### v2

@@ 1 Introduction
Specifically, Glyph consists of three main stages, namely, continual pre-training, LLM-driven rendering search, and post-training. In the continual pre-training stage, we render large-scale long-context text into diverse visual forms, enabling the VLM to transfer its long-context capability from text tokens to visual tokens. Since the text-to-image conversion directly determines the trade-off between context compression and model performance, devising an optimal configuration of the conversion is crucial for downstream performance. To this end, we design an LLM-driven genetic search to automatically explore rendering parameters (e.g., font size, layout, resolution) to maximize compression while preserving long-context ability. The resulting configuration is then applied in the post-training stage, where we perform supervised fine-tuning and reinforcement learning to further improve the model’s performance on visualized input. An auxiliary OCR task is applied to enhance the model’s ability to recognize textual content within images, thereby better aligning its visual and textual representations, yielding the final Glyph model.
@@
  • We demonstrate that Glyph can achieve 3-4× token compression for long text sequences while preserving performance, enabling substantial improvements in memory efficiency, training, and inference speed.
@@ 3.1 Overall Framework
As illustrated in Figure 2, Glyph consists of three tightly-coupled stages: (1) Continual Pre-Training, which teaches the VLM to understand and reason over rendered long texts with diverse visual styles; (2) LLM-Driven Rendering Search, automatically discovering the optimal rendering configuration for downstream tasks; and (3) Post-Training, including SFT and RL under the discovered configuration to further improve the model’s long-context capabilities. Together, these stages enable Glyph to achieve both high accuracy and significant gains in token compression, computational efficiency, and memory usage.
@@ 3.2 Task Definition
Scaling this token-based formulation to million-token contexts, however, imposes prohibitive memory and computation costs. To overcome these limitations, we reformulate the input representation through visual compression. Instead of directly feeding $\mathcal{C}$ as text tokens, we render it into a sequence of visual pages $\mathcal{V}=\{v_{1},\dots,v_{n}\}$, each containing the glyphs of multiple text segments. This allows the model to reason over a compressed but semantically equivalent input:
@@ 3.3 Continual Pre-Training
To enhance model robustness, better aligning long-text capability, we adapt diverse rendering configurations over a large amount of long-context text data. We also develop a series of rules to exclude the improper combination of rendering parameters, e.g., a smaller line height than font size. Moreover, with human prior, we define several style themes, including document_style, web_style, dark_mode, code_style, and artistic_pixel. These themes are designed to capture a wide range of document layouts and text styles, which can better exploit the knowledge that VLM has obtained in its pre-training stage.
@@
where $\mathcal{I}^{\!*}$ denotes an optional instruction (e.g., absent in interleaved language modeling tasks) and $\phi$ is initialized from the base VLM. This stage produces a model capable of understanding rendered text and handling long contexts, referred to as Glyph-Base.
@@ 3.5 Post-Training
Table 1: Performance comparison of Glyph with leading LLMs on LongBench (%). Our model achieves competitive results in the overall average score. Best results are bolded, and second-best are underlined. Refer to Table 10 for the rest of the results.
@@
Figure 3: Performance comparison of Glyph and the baseline across different context windows, demonstrating that Glyph achieves performance equivalent to longer contexts with substantially shorter context windows.
@@
  • Verifiable rewards from an external LLM judge, scoring based on the accuracy of the answer, which is a reference-based LLM-as-a-judge with the reference being the ground truth.
@@
A persistent challenge of visual compression is the faithful recovery of fine-grained text from rendered images. Throughout both SFT and RL, we therefore incorporate an auxiliary OCR alignment task that encourages the model to correctly read and reproduce low-level textual details. The form of the OCR task is the same as in the continual pre-training stage. In the RL stage, the reward for the OCR task is given by the Levenshtein distance.
@@ 4.1 Experimental Setup
To comprehensively evaluate the effectiveness of our method, we have conducted extensive experiments covering long-context understanding, efficiency, cross-modal generalization, and several ablations and analysis. Implementation details, descriptions of baselines and benchmarks are provided in Appendix B.
@@ 4.2.1 Results on LongBench & MRCR
Tables 1 and 2 summarize overall results. Glyph achieves performance on par with or surpassing state-of-the-art text-only LLMs of similar size, including Qwen3-8B and GLM-4-9B-Chat-1M, demonstrating that Glyph remains effective on long-context tasks with a large reduction in input tokens.
@@ 4.5 Ablation Study & Analysis
We compare three types of rendering configurations for SFT: (i) randomly sampled configuration from the pre-training sets, (ii) manually designed settings based on prior knowledge, and (iii) the configuration obtained from our search procedure. While all settings achieve comparable compression ratios, Table 5 shows that the searched configuration consistently outperforms the other two, both on average and across most individual tasks. This demonstrates the importance of systematic exploration for finding appropriate rendering strategies.
@@
To further examine the potential of our approach, we explore more aggressive compression settings. We apply a configuration with an effective $8\times$ compression ratio during post-training, and evaluate the resulting model on MRCR with sequence lengths extended from 128k to 1024k. The results (Table 7) show that Glyph successfully demonstrates the potential for $8\times$ effective context expansion, achieving performance on par with GLM-4-9B-Chat-1M and Qwen2.5-1M. This experiment highlights that our method can indeed be pushed to more extreme compression regimes while retaining performance, suggesting substantial headroom for extending usable context far beyond current limits, like a model that can deal with 4M, even 8M context tokens.
@@ 5 Conclusion
Despite the effectiveness of Glyph and its strong potential for broader applications, we want to discuss several limitations of the current work that are worth further exploration.
@@
Building upon the current study, several directions could further advance the proposed visual–text compression paradigm. First, rather than using a fixed rendering strategy, one promising avenue is to train adaptive rendering models that condition on the task type or user query, producing tailored visualizations that balance compression and performance. Second, enhancing the visual encoder’s capability for fine-grained text recognition and alignment with language representations could improve robustness and transferability across tasks. Third, improving the alignment between visual–text and purely textual models, for instance, through knowledge distillation or cross-modal supervision, could narrow the performance gap in generalization. Fourth, our approach could be extended to broader applications, such as agent memory systems capable of managing long-term conversations or agentic contexts, and tasks that can leverage structured visual layouts for reasoning and retrieval. From the perspective of context engineering, this method offers a new way to optimize how contextual information is represented and managed. With further advances along this line, future models could go beyond the current limits of context length, effectively scaling from 1M to 10M input tokens.
@@ Appendix A Rendering Parameters
Our rendering parameters are detailed in Table 8. The best result, obtained through our LLM-driven genetic search, is presented in Figure 6, which shows the detailed configuration and its corresponding rendering.
@@
dpi Mixture of sets: lowest (45–59), low (60–71), medium (72–119), normal ({72,80,96,100,110,120,144,150,300}), high (over 300); favor normal/medium with small probability spikes to extremes.
@@
Table 10: The rest of the results on LongBench benchmark (%), which encompasses Single-Document QA, Multi-Document QA, Summarization, Few-shot Learning, and Synthetic task.
@@ Appendix B Implementation Details
For the rendering search, we run for 5 times with 200 steps in each round, to find the optimal configuration that maximizes the compression ratio while maintaining good performance.
After this, we conduct further SFT and RL training. For SFT, we train for 1.5k steps with a batch size of 32. The Adam optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.95$) is used with cosine decay and 160 warm-up steps, where the learning rate decays from 5e-6 to 2e-6. For reinforcement learning, we adopt the GRPO algorithm. Each training group samples 16 candidate responses, and degenerate samples with all-zero or all-one rewards are discarded. We apply the clip-higher trick from DAPO Yu et al. (2025b) with $\epsilon_{l}$ being 0.2 and $\epsilon_{h}$ being 0.28. Training runs for 500 iterations with a batch size of 32. We also use the Adam optimizer with a constant learning rate of 1e-6.
@@
Our method relies on a strong VLM to process long-context tasks. Considering the impressive performance of GLM-4.1V-9B, especially in OCR and long document tasks, we have chosen GLM-4.1V-9B-Base as our backbone model.
@@
To conduct a comprehensive analysis of the long-context performance, we have adopted three popular benchmarks, including LongBench, MRCR, and Ruler. LongBench consists of 21 datasets in total in 6 categories, covering diverse long-context tasks. MRCR is a task proposed by Vodrahalli et al. (2024). We use the OpenAI version, which consists of multi-turn conversations about writing, asking models to recall one of the contexts in dialogue history. Ruler is a widely used synthetic benchmark with 11 NIAH tasks. To validate cross-model benefits, we choose the MMLongBench-Doc, which involves 130 lengthy PDF with diverse layout and images, and 1062 questions.
@@
For training, we focus on SFT since RL involves rollout time, making it difficult to compare fairly. Moreover, running RL at very long context lengths (e.g., 64k or beyond) requires excessive memory resources. This again highlights the advantage of our approach: through compression, we can conduct RL training at 32k context length while effectively covering over 100k tokens of raw text input, where RL is prohibitively difficult for LLMs due to memory and computation demands. For SFT, we measure per-sample training time under the same number of training data using $\times8$0G H100 GPUs. For inference, we deploy both models on a single 80G H100 and measure efficiency along two axes: (i) prefill latency at batch size 1, and (ii) per-sample inference time at the maximum feasible batch size with output length set to 256 tokens. We omit the KV cache testing, because KV cache scales linearly with the sequence length, and compression translates almost directly into savings of about 67% memory usage.